# Modelling Global Trade with Optimal Transport

### v2

@@ Results
In January 2018, the Trump administration started imposing import tariffs on goods primarily from China. In response, the Chinese government increased tariffs on a variety of products, including agricultural imports. The largest agricultural export from the US to China, soya beans, were hit with a 25% import tariff [30]. Meanwhile, political tensions between China and Australia caused Beijing to introduce high anti-dumping tariffs on Australian exports such as barley (80.5%) and wine (206%), starting in 2020 [31]. Wine trade had previously been tariff-free since the signing of ChAFTA in 2015. Our analysis provides an estimate of the change in the ease of trading these measures induced (figs. LABEL:fig:China_tradeB and LABEL:fig:Barley_trade). Australian beef, wine and barley imports all experience large increases in cost, following the implosion of trade volumes. Australia was able to divert some of its excess barley supply to Saudi Arabia, which saw a decrease in trade barriers of over 0.5 points between 2019 and 2022 (fig. LABEL:fig:Barley_tradeC). Trade volumes to Vietnam also increased from 200,000 to 800,000 metric tons, though trade costs remained approximately constant. Meanwhile, after 2020 China doubled its barley imports from Canada and France. We found that import barriers from both countries were reduced slightly in 2021, though they increased again in the year after.
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] : Figure 7Comparison with gravity model. A–B Comparison plot of the OT and gravity estimates ($x$-axis) versus the true data ($y$-axis) on two selected commodities. Also shown is a linear fit (dotted line), its estimated slope $m$, the Pearson coefficient of the fit $r$, and the line $y=x$ (solid line). See fig. LABEL:fig:Gravity_comparisons_all in the appendix for an overview of all commodities. C–D Comparison of the RMSE accuracies of the estimated transport volumes of the OT approach (blue) and the gravity model (orange). Values are averaged over all countries and years, with the errorbars showing one standard deviation from the mean (triangular marker). Also shown are the median values (diamond markers). Shown are the RMSE (left) and the RMSE in units of the standard deviation on the true data (right).
@@
In 2016, the United Kingdom voted to leave the European Union, officially exiting the common market and customs union on December 31, 2020. This case study examines the impact of Brexit on British import patterns by comparing vegetable and wine imports from mainland Europe to both the United Kingdom and the Republic of Ireland (ROI), which remains part of the Eurozone and the common market. While both island nations naturally source the majority of their fresh produce from mainland Europe, their trading patterns have evolved in markedly different ways. Imports of lettuce from Europe generally fell for the UK, accompanied by a rise in import cost: –44% trade volume and +0.11 in import costs from the Netherlands, the largest exporter of lettuce and chicory to the UK, as well as a –21% drop in trade from Spain, though with no change in import cost. Ireland increased its imports of lettuce and other greens from the Netherlands, Spain, Italy, and Portugal, accompanied by a general decrease in trading costs. In the case of the Netherlands, Ireland saw a consistent reduction in vegetable trade costs, unlike the United Kingdom. It is interesting to note that the United Kingdom significantly increased its imports of vegetables from Morocco, accompanied by a precipitous drop in trade costs, indicating a facilitation of trade between the two countries in the wake of Brexit. This is not true for the ROI: though it increased its imports of Moroccan tomatoes by nearly twice as much as the UK, Irish trade costs still fell by less than for the UK.
A more clear-cut trend emerges in the wine trade (fig. LABEL:fig:BrexitB): here, the UK was consistently affected more negatively than the Republic of Ireland: British import costs from all eight countries considered rose by considerably more than those of the ROI. An 11% drop in Spanish wine import was accompanied by a 0.09 point increase in trading costs, while a –8% change in Irish imports was nonetheless accompanied by a 0.01 point decrease in costs. Portugese wine imports to the UK rose by 24%, notwithstanding a 0.07 increase in trade costs. The picture is unchanged for South African, Australian, and New Zealand imports. The EU maintains free trade or regulatory agreements removing wine import duties with the former two [32, 33]. When the UK left the European Union, wine from Australia entered at the UK Global Tariff rate, which in mid-2023 was eliminated under the Australia-United Kingdom FTA [34]. South African wines, by contrast, continued to be imported to the UK tariff-free post-Brexit [35]. Yet here too, the United Kingdom’s 5% decrease in imports was driven by a 0.07 increase in trading costs; Ireland, by contrast, imported an estimated 61% less wine, driven by a comparable 0.1 point increase in trading costs.
@@
Lastly, we compare the performance of our method with a standard gravity model [11, 36, 37, 38], as specified in equations (1)–(2). The covariates include geographic distance, shared borders, colonial ties, common language, regional trade agreements, tariffs, and importer/exporter fixed effects to account for multilateral resistance (see Supplementary Information for details). We estimate the coefficients using Poisson Pseudo Maximum Likelihood estimation and compare the accuracy of the estimated transport plans $\mathbf{T}$. Figures LABEL:fig:Gravity_compA–B show scatter plots of the OT (blue) and gravity (orange) estimates against the FAO data. For all commodities studied, a linear fit through the OT estimates yields a near-perfect slope of $m=1$ with a Pearson coefficient close to 1, perfectly fitting the tail end of the distribution. In contrast, the gravity model’s performance is much more volatile, with linear fits ranging from a Pearson coefficient of between 0.998 (best) to 0.747 (worst) (see also Supplementary Information). Due to model misspecification, the fits to the tails of the distributions are generally significantly poorer. Consequently, figure LABEL:fig:Gravity_compC shows that the OT approach significantly outperforms the gravity model in terms of RMSE, often by two to three orders of magnitude. Figure LABEL:fig:Gravity_compD illustrates that OT estimates typically fall within one standard deviation of the data uncertainty, whereas gravity estimates tend to range from one to two, at times even three to four standard deviations. The gravity model also exhibits much higher variance in accuracy compared to OT.
@@ Discussion
This paper introduces a novel and versatile approach for identifying the drivers and barriers of global commodity trades. Using optimal transport theory, we are able to obtain a cost structure that is more expressive than a covariate-based gravity approach. Our estimates are thus orders of magnitude more accurate than the current state of the art, while providing consistent accuracy across datasets. The optimal transport approach models trade networks as a dynamical, interconnected system, allowing to capture complex rearrangements and network response dynamics to e.g. trade wars, conflicts, or shifts in political relations. By contrast, the covariate-based gravity approach fits each node in the network individually without taking interaction effects into account. Though the current work looks only at global agrifood markets, the methodology proposed is general and applicable to commodity flows, financial markets, or banking networks [39]. Beyond economics, the optimal transport approach also relates e.g. to global migration flows, which can be estimated from stock data [40, 41].
Method
@@
In OT one wishes to find the optimal flow of mass from a source distribution to a target distribution, while minimising an overall transport cost. This abstract problem has a wide range of applications in economics, logistics, image restoration, transport systems, or urban structure [21, 42, 43].
Consider an $m$-dimensional space $X$, an $n$-dimensional space $Y$, and furthermore a non-negative measure $\mathbf{C}$ on $X\times Y$. The entries of $\mathbf{C}$ correspond to the cost of transporting mass from one location in $X$ to a target in $Y$. Given two probability measures $\bm{\mu}\in P(X)$ and $\bm{\nu}\in P(Y)$ (the supply and demand), the OT problem consists in finding a transport plan $\mathbf{T}$ minimising the overall cost eq. (3). The transport plan $\mathbf{T}$ must also satisfy the marginal constraints
@@
In practice one usually solves the entropy regularised OT formulation, which can be solved much more efficiently [44]; here, an additional term is added to the objective:
$\min_{\mathbf{T}}\sum_{ij}C_{ij}T_{ij}+\epsilon\sum_{ij}T_{ij}\left(\log T_{ij}-1\right),$ (10)
where $\epsilon>0$ is a positive regularisation parameter. This regularisation prevents monopolisation, i.e. demand being supplied from only a few sources.
@@
$+\epsilon\sum_{ij}T_{ij}\left(\log T_{ij}-1\right),$ (11)
with $\bm{\lambda}\in\mathbb{R}^{m}$ and $\bm{\eta}\in\mathbb{R}^{n}$ Lagrangian multipliers. Minimising $\mathcal{L}$ with respect to $\mathbf{T}$ gives the solution
$T_{ij}=e^{-\lambda_{i}/\epsilon}e^{-C_{ij}/\epsilon}e^{-\eta_{j}/\epsilon}$ (12)
@@
$\mathbf{T}=\bm{\Pi}e^{-\mathbf{C}/\epsilon}\bm{\Omega},$ (13)
where $\bm{\Pi}=\mathrm{diag}(e^{-\lambda_{1}/\epsilon},\cdots,e^{-\lambda_{m}/\epsilon})\in\mathbb{R}^{m\times m}$ and $\bm{\Omega}=\mathrm{diag}(e^{-\eta_{1}/\epsilon},\cdots,e^{-\eta_{n}/\epsilon})\in\mathbb{R}^{n\times n}$ are diagonal matrices of Lagrangian multipliers.
Finding $\bm{\Pi}$ and $\bm{\Omega}$ is achieved through an iterative procedure that is variously called Iterative Proportionate Procedural Fitting (IPFP), RAS, or Sinkhorn’s algorithm [45, 44, 46]. Define $\mathbf{M}=e^{-\mathbf{C}/\epsilon}$; then, given an initial guess $\bm{\Pi}^{0}$, we update $\bm{\Omega}$ to satisfy the first marginal constraint eq. (9)
$\bm{\Omega}\mathbf{M}^{\top}\bm{\Pi}^{0}=\bm{\nu}.$ (14)
@@
2:      $\mathbf{M}$ (Exponential of cost matrix) $\bm{\mu},\bm{\nu}$ (marginals)
@@
5:     $\bm{\Omega}^{i+1}\leftarrow\dfrac{\bm{\nu}}{\mathbf{M}\bm{\Pi}^{i}}$
6:     $\bm{\Pi}^{i+1}\leftarrow\dfrac{\bm{\mu}}{\mathbf{M}^{\top}\bm{\Omega}^{i+1}}$
@@
Under certain conditions, convergence of the algorithm to a unique solution is guaranteed [47, 48]. Note the important fact that the solution $\mathbf{T}$ is invariant under scaling of the cost matrix, since the Lagrangian multipliers absorb the scaling; the transport plan thus does not depend on absolute cost values, only on their relative proportions.
The classic OT problem eq. (6) can be interpreted as the central’s planners problem of finding the optimal assignment/matching of supplies and demands. The dual OT problem is given by
$\max_{\bm{f},\bm{g}}\langle\bm{f},\bm{\mu}\rangle+\langle\bm{g},\bm{\nu}\rangle$ (17)
such that $\bm{f}\in\mathbb{R}^{n}$ and $\bm{g}\in\mathbb{R}^{m}$ satisfy
@@
where $\bm{f}\oplus\bm{g}=\bm{f}\bm{1}^{T}_{m}+\bm{g}\bm{1}^{T}_{n}$. Here $\bm{f}$ and $\bm{g}$ can be interpreted as the minimal cost of picking up and dropping off a good at locations respectively. The central planner problem of finding the best plan $\mathbf{T}$ is therefore split into determining the optimal cost of collecting and delivering goods. The constraint (18) ensures optimality. If $f_{i}+g_{j}>C_{ij}$, that is the cost of picking up a good at location $i$ and dropping it off at location $j$ is larger than the transportation cost, it can not be optimal.
@@
To infer the cost matrix function $\mathbf{C}(t)$ from a dataset of transport plan observations $\mathbf{T}(t)$, we build on the neural parameter estimation method first introduced in [23] and subsequently expanded upon [24]. We wish to train a neural network $u$ to solve the inverse OT problem $\mathbf{C}(t)=u(\mathbf{T}(t))$. We do so by constructing a loss function that differentiates through the optimal transport equations, i.e.
$J=\|\mathbf{\hat{T}}(\mathbf{\hat{C}})-\mathbf{T}\|_{2}^{2}+\|\sum_{j}C_{ij}-1\|_{1}.$ (19)
Here, $\mathbf{\hat{T}}(\mathbf{\hat{C}})$ is the estimated transport plan obtained by solving Sinkhorn’s algorithm alg. [1]. The second summand together with the restriction that $0\leq C_{ij}\leq 1$ is necessary to fix the cost matrix since the OT problem is in general invariant under affine transformations of $\mathbf{C}$. Using this approach, the quality of the prediction does not depend on the number of datasets used to train the neural network, since we are not performing regression, but rather fitting a mathematical model (or a set of parameters) to data. The data is processed in batches, and a gradient descent step performed on the neural network parameters after each batch. The loss is only calculated for links with trade flow $>0$.
@@
As mentioned, the FAO dataset contains two values for each entry $T_{ij}$: one reported by the exporter, and one by the importer. Let $\mathbf{T}^{\text{E}}$ by the transport plan where all entries are those reported by the exporters, and $\mathbf{T}^{\text{I}}$ those where all are reported by the importers. The training data—i.e., the data we use to train the function $u$—consists of only these two transport plans for each year: $\{\mathbf{T}^{\text{E}}(t),\mathbf{T}^{\text{I}}(t)\}$, giving a total training set size of $2\times L$, where $L=23$ are the number of observation points. A hyperparameter sweep showed that using a deep neural network with 5 layers, 60 nodes per layer, and hyperbolic tangent activation functions on all layers but the last, where we use a sigmoid, gives best results. Using a sigmoid activation function on the last layer ensures $0\leq C_{ij}\leq 1$. We use the Adam optimizer [49] to train the neural network. We pool all FAO trade matrices to only contain those countries that account for 99% of import and export volumes, subsuming all other countries in an ‘Other’ category (thereby ensuring that no flow is lost). Missing entries in the training data are masked and do not contribute to the loss function.
Uncertainty quantification on the estimated cost matrix $\hat{\mathbf{C}}$ is obtained by passing random samples of $\mathbf{T}$ through the trained neural network $u$. These samples are obtained by selecting either $T_{ij}^{\text{E}}$ or $T_{ij}^{\text{I}}$ uniformly at random for each entry of the transport plan, and passing this sample through the neural network. Repeating this $n$ times gives $n$ samples of $\mathbf{C}$, and inserting each estimate of $\mathbf{C}$ into Sinkhorn’s algorithm gives $n$ estimated transport plans $\mathbf{\hat{T}}$. We generate $n=1000$ samples for each year.

### v3

@@ Results
In January 2018, the first Trump administration started imposing import tariffs on goods primarily from China. In response, the Chinese government increased tariffs on a variety of products, including agricultural imports. The largest agricultural export from the US to China, soya beans, were hit with a 25% import tariff [Biesheuvel_2018]. Meanwhile, political tensions between China and Australia caused Beijing to introduce high anti-dumping tariffs on Australian exports such as barley (80.5%) and wine (206%), starting in 2020 [Sullivan_2022]. Wine trade had previously been tariff-free since the signing of ChAFTA in 2015. Our analysis provides an estimate of the change in the ease of trading these measures induced (figs. LABEL:fig:China_tradeB and LABEL:fig:Barley_trade). Australian beef, wine and barley imports all experienced large increases in cost, following the implosion of trade volumes. Australia was able to divert some of its excess barley supply to Saudi Arabia, which saw a decrease in trade costs of over 0.8 points between 2019 and 2022 (fig. LABEL:fig:Barley_tradeC). Trade volumes to Vietnam also increased from 200,000 to 800,000 metric tons, though trade costs remained approximately constant. Meanwhile, after 2020 China doubled its barley imports from Canada and France. We found that import barriers from both countries were reduced only slightly in 2021 and rebounded the following year.
@@
A Estimates vs FAO data, cucumbers B Estimates vs FAO data, barley [Uncaptioned image] [Uncaptioned image] C Estimation accuracy (RMSE) D Estimation accuracy (RMSE in standard deviations) [Uncaptioned image] : Figure 6Comparison with gravity model. A–B Comparison plot of the OT and gravity estimates ($x$-axis) versus the true data ($y$-axis) on two selected commodities. Also shown is a linear fit (dotted line), its estimated slope $m$, the Pearson coefficient of the fit $r$, and the line $y=x$ (solid line). See fig. LABEL:fig:SI:Gravity_comparisons_all in the appendix for an overview of all commodities. C–D Comparison of the RMSE accuracies of the estimated transport volumes of the OT approach (blue) and the gravity model (orange). Values are averaged over all countries and years, with the errorbars showing one standard deviation from the mean (triangular marker). Also shown are the median values (diamond markers). Shown are the RMSE (C) and the RMSE in units of the standard deviation on the true data (D).
@@
In 2016, the United Kingdom voted to leave the European Union, officially exiting the common market and customs union on December 31, 2020. This case study examines the impact of Brexit on British import patterns by comparing vegetable and wine imports from mainland Europe to both the United Kingdom and the Republic of Ireland (ROI), which remains part of the Eurozone and the common market. While both island nations naturally source the majority of their fresh produce from mainland Europe, their trading patterns have evolved in markedly different ways (fig. LABEL:fig:BrexitA). Imports of lettuce from Europe generally fell for the UK, accompanied by a rise in import cost: 25% decrease in trade volume and +0.04 in import costs from the Netherlands, the largest exporter of lettuce and chicory to the UK, as well as a –28% drop in trade from Spain, though with no change in import cost. Ireland increased its imports of lettuce and other greens from the Netherlands, Spain, and Italy, accompanied by a general decrease in trading costs for those products. The ROI’s imports of greenery from Portugal fell by 53%, with no change in trade barriers; for the UK, a 71% fall in trade volumes accompanied a 0.12 point increase in trade cost. In the case of the Netherlands, Ireland saw a consistent reduction in vegetable trade costs, unlike the United Kingdom. It is interesting to note that the United Kingdom significantly increased its imports of vegetables from Morocco, accompanied by a precipitous drop in trade costs, indicating a facilitation of trade between the two countries in the wake of Brexit. This is not true for the ROI: though it increased its imports of Moroccan tomatoes and cucumbers, Irish trade costs remained mostly unchanged.
A more clear-cut trend emerges in the wine trade (fig. LABEL:fig:BrexitB): here, the UK was consistently affected more negatively than the Republic of Ireland: British import costs from all eight countries considered rose by considerably more than those of the ROI. A 9% drop in Spanish wine import by the UK was accompanied by a 7% increase in trading costs, while a 13% drop in Irish imports only meant a 5% increase in costs. Portugese wine imports to the UK rose by 16%, notwithstanding a 0.05 increase in trade costs. A similar pattern holds for South African, Australian, and New Zealand imports. The EU maintains free trade or regulatory agreements removing wine import duties with the former two [EU-SADC, WFA_2016]. When the UK left the European Union, wine from Australia entered the UK Global Tariff rate, which in mid-2023 was eliminated under the Australia-United Kingdom FTA [DAFF_2020]. South African wines, by contrast, continued to be imported to the UK tariff-free post-Brexit [VINPRO_2020]. Yet here too, the United Kingdom’s 13% decrease in imports was driven by a 0.07 increase in trading costs; Ireland, by contrast, imported 19% less wine, driven by a comparable 0.07 point increase in trading costs.
@@
Lastly, we compare the performance of our method with a traditional gravity model [Yotov_2017, CEPII, UNCTAD_TRAINS, WTO_Stats], as specified in equations (1)–(2). The covariates include geographic distance, shared borders, colonial ties, common language, regional trade agreements, tariffs, and importer/exporter fixed effects. We estimate the coefficients using Poisson Pseudo Maximum Likelihood estimation [silva2006log, Correia2020] and compare the accuracy of the estimated transport plans $\bm{T}$. Figures LABEL:fig:Gravity_compAB show scatter plots of the OT (blue) and the gravity (orange) estimates against the FAO data. For all commodities studied, a linear fit through the OT estimates yields a near-perfect slope of $m=1$ with a Pearson coefficient close to $1$, particularly providing an exact fit for the upper tail of the trade value distribution. In contrast, the gravity model’s performance is much more volatile, with linear fits ranging from a Pearson coefficient of between 0.975 (best) to 0.699 (worst) (see also SI). Due to model misspecification, the fits to the tails of the distributions are generally significantly poorer. Consequently, figure LABEL:fig:Gravity_compC shows that the OT approach significantly outperforms the gravity model in terms of RMSE, often by an order of magnitude. Figure LABEL:fig:Gravity_compD illustrates that OT estimates typically fall within one standard deviation of the data uncertainty, whereas gravity estimates tend to range from one to two, at times even three to four standard deviations. The gravity model also exhibits much higher variance in accuracy compared to OT. We further an investigate an alternative specification by replacing time-varying country-level and time-invariant pair regressors with exporter-time, importer-time, and exporter-importer fixed effects in a three-way gravity model [Breinlich2021], capturing multilateral resistance terms more effectively (see eq. (21) in the SI for details). As this formulation incorporates higher degrees of freedom in the three-way specification, its performance naturally converges toward that of OT, which more accurately fits high-value trade flows. Incorporating time-varying bilateral trade regressors would likely further narrow this gap, reinforcing the validity of the OT approach. Notably, optimal transport provides trade cost estimates even without such regressors—many of which are difficult to quantify.
@@ Discussion
This paper introduces a novel and versatile approach for identifying the drivers and barriers of global commodity trades. Using optimal transport theory, we are able to obtain a cost structure that is more expressive than a covariate-based gravity approach. Our estimates are thus orders of magnitude more accurate than traditional gravity models, matching the performance of high-dimensional fixed-effects specifications while maintaining consistent accuracy across datasets. The optimal transport approach models trade networks as a dynamical, interconnected system, allowing to capture complex rearrangements and network response dynamics to e.g. trade wars, conflicts, or shifts in political relations. Though the current work looks only at global agrifood markets, the methodology proposed is general and applicable to commodity flows, financial markets, or banking networks [Giesecke2019]. Beyond economics, the optimal transport approach also relates e.g. to global migration flows, which can be estimated from migrant stock data [Willekens_1999, Azose_2018, Gaskin_Abel_2025]. Future work could explore correlations between related commodities within this framework, as well as develop hybrid models that combine observed covariates with data-driven residuals via semi-structured OT costs. Another promising direction is counterfactual analysis using the conditional equilibrium framework of [Yotov_2017]. Extending the model to a full general equilibrium setting with endogenous production and pricing is also a natural next step.
Methods
@@
In OT one wishes to find the optimal flow of mass from a source distribution to a target distribution, while minimising an overall transport cost. This abstract problem has a wide range of applications in economics, logistics, image restoration, transport systems, or urban structure [Galichon_2016, Santambrogio_2015, Peyre_2019].
Consider an $m$-dimensional space $X$, an $n$-dimensional space $Y$, and $\bm{C}$ a measure on $X\times Y$. The entries of $\bm{C}$ correspond to the cost of transporting mass from one location in $X$ to a target in $Y$. Given two probability measures $\bm{\mu}\in P(X)$ and $\bm{\nu}\in P(Y)$ (the supply and demand), the OT problem consists in finding a transport plan $\bm{T}$ minimising the overall cost eq. (3). The transport plan $\bm{T}$ must also satisfy the marginal constraints
@@
In practice one usually solves the entropy regularised OT formulation, which can be solved much more efficiently [Cuturi_2013]; here, an additional term is added to the objective:
$\min_{\bm{T}}\sum_{ij}C_{ij}T_{ij}+\varepsilon\sum_{ij}T_{ij}\left(\log T_{ij}-1\right),$ (10)
where $\varepsilon>0$ is a positive regularisation parameter. This regularisation prevents monopolisation, i.e. demand being supplied from only a few sources.
@@
$+\varepsilon\sum_{ij}T_{ij}\left(\log T_{ij}-1\right),$ (11)
with $\bm{\lambda}\in\mathbb{R}^{m}$ and $\bm{\eta}\in\mathbb{R}^{n}$ Lagrangian multipliers. Minimising $\mathcal{L}$ with respect to $\bm{T}$ gives the solution
$T_{ij}=e^{-\lambda_{i}/\varepsilon}e^{-C_{ij}/\varepsilon}e^{-\eta_{j}/\varepsilon}$ (12)
@@
$\bm{T}=\bm{\Pi}e^{-\bm{C}/\varepsilon}\bm{\Omega},$ (13)
where $\bm{\Pi}=\mathrm{diag}(e^{-\lambda_{1}/\varepsilon},\cdots,e^{-\lambda_{m}/\varepsilon})\in\mathbb{R}^{m\times m}$ and $\bm{\Omega}=\mathrm{diag}(e^{-\eta_{1}/\varepsilon},\cdots,e^{-\eta_{n}/\varepsilon})\in\mathbb{R}^{n\times n}$ are diagonal matrices of Lagrangian multipliers.
Finding $\bm{\Pi}$ and $\bm{\Omega}$ is achieved through an iterative scaling procedure that is variously called Iterative Proportional Fitting (IPF), RAS, or Sinkhorn’s algorithm [Deming_1940, Cuturi_2013, Idel_2016]. Define $\bm{M}=e^{-\bm{C}/\varepsilon}$; then, given an initial guess $\bm{\Pi}^{0}$, we update $\bm{\Omega}$ to satisfy the first marginal constraint eq. (9)
$\bm{\Omega}\bm{M}^{\top}\bm{\Pi}^{0}=\bm{\nu}.$ (14)
@@
2:   $\bm{M}$ (Exponential of cost matrix) $\bm{\mu},\bm{\nu}$ (marginals)
@@
5:  $\bm{\Omega}^{i+1}\leftarrow\dfrac{\bm{\nu}}{\bm{M}\bm{\Pi}^{i}}$
6:  $\bm{\Pi}^{i+1}\leftarrow\dfrac{\bm{\mu}}{\bm{M}^{\top}\bm{\Omega}^{i+1}}$
@@
Under certain conditions, convergence of the algorithm to a unique solution is guaranteed [Sinkhorn_1964, Rueschendorf_1995].
The classic OT problem eqs. (3)–(4) can be interpreted as the central planner’s problem of finding the optimal assignment or matching of supply and demand. The entropy-regularised OT problem can be viewed as a similar optimal assignment problem, but subject to uncertainty and/or randomisation. The dual problem to entropy-regularised OT, i.e. minimising (5) subject to (3), is given by
$\max_{\bm{f},\bm{g}}\langle\bm{f},\bm{\mu}\rangle+\langle\bm{g},\bm{\nu}\rangle{-\varepsilon\sum_{i,j}e^{\frac{f_{j}+g_{i}-C_{ij}}{\varepsilon}}}.$ (17)
In the limit $\varepsilon\rightarrow 0$ the last term ensures that the dual potentials $\bm{f}$ and $\bm{g}$ satisfy
@@
where $\bm{f}\oplus\bm{g}=\bm{f}\bm{1}^{T}_{m}+\bm{g}\bm{1}^{T}_{n}$. Condition (18) corresponds to the admissibility condition of the dual non-regularised OT problem. In this context $\bm{f}$ and $\bm{g}$ can be interpreted as the minimal cost of picking up and dropping off a good at locations respectively. The problem of finding the optimal plan $\bm{T}$ is thus split into determining the optimal cost of collecting and delivering goods. The constraint (18) ensures optimality: if $f_{i}+g_{j}>C_{ij}$, that is, the cost of picking up a good at location $i$ and dropping it off at location $j$ is larger than the transportation cost, it cannot be optimal.
@@
In inverse optimal transport one wishes to infer the underlying cost $\bm{C}$ from (partial) observations of transport plans $\bm{T}$, which are usually solutions to entropy regularised OT problems. Rewriting eq. (13), we have
$\bm{C}=\varepsilon\left(\log\bm{\Pi}+\log\bm{\Omega}^{\top}-\log\bm{T}\right);$
thus $\bm{C}$ is only determined up to an additive decomposition into row and column vectors, with any transformation of the kind
$C_{ij}\mapsto C_{ij}+\alpha_{i}+\beta_{j}$
leaving the transport plan $\bm{T}$ invariant, since the transformation can be absorbed by the scaling vectors. To constrain the problem, we can bound the cost $C_{ij}\in[0,C_{max}]$, and demand that the inverse of 0 should be $C_{max}$, i.e.
$u:\bm{T}\mapsto\bm{C},\,u(0)=C_{max}.$
This is a natural restriction, since it implies that the cost on edges with zero transport flow should be maximal. If this maximum is attained in every row and every column of the cost matrix, the row- and column-shifts $\alpha_{i}$ and $\beta_{j}$ must satisfy
$C_{max}=\max_{j}C_{ij}\overset{!}{=}\max_{j}(C_{ij}+\alpha_{i})\leq C_{max}\ \forall\ i\Rightarrow\alpha_{i}=0.$
(and similarly $\beta_{j}$). Thus, under these conditions the cost matrix is uniquely determined (see also fig. LABEL:fig:SI:validation in the SI).
@@
To infer the cost matrix function $\bm{C}(t)$ from a dataset of transport plan observations $\bm{T}(t)$, we build on the neural parameter estimation method first introduced in [Gaskin_2023] and subsequently expanded upon [Gaskin_2024]. We wish to train a neural network $u$ to solve the inverse OT problem $\bm{C}(t)=u(\bm{T}(t))$. We do so by constructing a loss function based on the optimal transport equations, i.e.
$J=\|\bm{\hat{T}}(\bm{\hat{C}})-\bm{T}\|_{2}^{2}+\sum_{(i,j)\in\mathcal{S}}(C_{ij}-C_{max})^{2}.$ (19)
[…]
$\mathcal{S}:=\{(i,j)\ |\ T_{ij}=0\}$
[…]
@@
As mentioned, the FAO dataset contains two values for each entry $T_{ij}$: one reported by the exporter, and one by the importer. Let $\bm{T}^{\text{E}}$ be the transport plan where all entries are those reported by the exporters, and $\bm{T}^{\text{I}}$ those where all are reported by the importers. The training data—i.e., the data we use to train the function uu—consists of only these two transport plans for each year: $\{\bm{T}^{\text{E}}(t),\bm{T}^{\text{I}}(t)\}$, giving a total training set size of $2\times L$, where $L=23$ are the number of observation points. A hyperparameter sweep showed that using a deep neural network with 5 layers, 60 nodes per layer, and hyperbolic tangent activation functions on all layers but the last, where we use a sigmoid, gives best results. Using a sigmoid activation function on the last layer ensures $0\leq C_{ij}\leq 1=C_{max}$. We use the Adam optimizer [Kingma_2014] to train the neural network. We pool all FAO trade matrices to only contain those countries that account for 99% of import and export volumes, subsuming all other countries in an ‘Other’ category (thereby ensuring that no flow is lost). Entries for which neither the importer nor the exporter have reported a value are assumed to be zero, and we constrain the cost matrix to be maximal on these entries. Entries for which only the exporter or the importer have reported a value (but not the other) are presumed missing in the respective table, and are masked in the loss function. With this approach, on average about 20% of entries are masked in the transport plan (see fig. LABEL:fig:SI:masking in the SI). Entries for which all reported values are missing populate the zero-flow edge set $\mathcal{S}$.
Uncertainty quantification
Uncertainty on the estimates stems from two sources: one, the degree to which the minimizer of the inverse problem (19) is ill-defined (i.e. the number of possible cost matrices that all fit the problem equally well), and two, the uncertainty on the transport plans themselves. To address the first, we use an ensemble training approach [Gaskin_2023, Gaskin_2024] and train a family $\{u_{k}\}$ of 10 neural networks for each commodity in parallel. Even though the inverse problem is theoretically well-posed when the forward problem (10) admits a solution, in the case of the FAO data we are solving a minimisation problem. To incorporate the uncertainty on the transport plans, we pass random samples of $\bm{T}$ through each of the trained neural networks $u_{k}$. These samples are obtained by selecting either $T_{ij}^{\text{E}}$ or $T_{ij}^{\text{I}}$ uniformly at random for each entry of the transport plan, and passing this sample through each neural network. Repeating this $n$ times gives $n$ samples of $\bm{C}$, and inserting each estimate of $\bm{C}$ into Sinkhorn’s algorithm gives $n$ estimated transport plans $\bm{\hat{T}}$. We generate a total of $n=100$ samples for each year and neural network (see also fig. LABEL:fig:SI:UQ in the SI). The uncertainty estimates obtained by our method then provide an indication of how strongly a given set of trade flows informs the underlying cost.